# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 4 Adaptive Sampling Methods
The adaptive batch schemes proposed in this paper are based on a family of adaptive sampling methods for stochastic optimization problems. Although such methods are extended to problems with different structures (see Section 2.2 for details), here we focus on the simplest unconstrained setting.
@@ 4.1 Norm Test
Byrd et al. (2012) proposed an adaptive sampling method, called the norm test or the norm condition, that was derived according to the following fundamental fact in stochastic optimization. If $f(\cdot;\xi)$ is a convex and continuously differentiable function for any $\xi\in\mathcal{Z}$, then $-\nabla F_{\mathcal{B}}(x)$ is a descent direction for $F$ at $x\in\mathbb{R}^{d}$ if there exists $\eta\in\left[0,1\right[$ such that
@@
See Appendix B for its proof. When the sample size $n$ is large, $\nabla F$ is not available, but $\delta_{\mathcal{B}}(x)$ can be approximated by
@@
Therefore, as the sample size $n\to+\infty$, condition (3) can be approximated as
$\frac{\mathrm{Var}_{i\in\mathcal{B}}(\nabla f(x;\xi_{i}))}{b}\leqslant\eta^{2}\|\nabla F_{\mathcal{B}}(x)\|^{2},$
which also corresponds to the case of sampling with replacement. In a practical implementation, the dynamic sample gradient algorithm with the norm test (Byrd et al., 2012) performs the following SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with step size $\alpha_{k}\in\mathbb{R}_{++}$, and then checks the following (approximate) norm test condition:
@@
meaning that the condition (3) holds in expectation. Here we abbreviate the conditional expectation with respect to $x_{k}$ (i.e., the sigma algebra $\mathcal{F}_{k}\coloneqq\sigma(\{x_{1},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$. Should the condition (5) not be met, the subsequent batch size is determined by the formula:
$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))}{\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right\rceil,$
and the batch $\mathcal{B}_{k+1}$ expanded accordingly. The selection of the constant $\eta$ in practice influences the probability of obtaining a descent direction, thereby affecting the probability of increasing the batch size.
@@ 4.2 Inner Product Test
Bollapragada et al. (2018a) noted that the norm test often leads to a rapid increase in batch size, which could negate the benefits of adaptive sampling, and proposed the inner product test as an alternative. This test further controls the variance of the term $\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle$ to ensure that $-\nabla F_{\mathcal{B}_{k}}(x_{k})$ is a descent direction with high probability. The test selects a large enough batch size $b_{k}$ to meet the (exact variance) inner product test condition, which is satisfied when there is a ϑ++italic-$\vartheta\in\mathbb{R}_{++}$ ensuring that:
@@
where the variance on the left hand side can be evaluated by (4). To establish a convergence rate, it is crucial to eliminate the possibility that $\nabla F_{\mathcal{B}_{k}}(x_{k})$ and $\nabla F(x_{k})$ are (near) orthogonal. Motivated by this need for theoretical guarantees, Bollapragada et al. (2018a) also introduce the following condition of the (exact variance) orthogonality test: there exists $\nu\in\mathbb{R}_{++}$ such that
@@
The above exact test also takes an approximate form for implementation purpose, given by
@@
Note that both (8) and (10) have to be satisfied, which collectively leads to the (approximate) augmented inner product test. If any of the conditions is violated, the batch size will be updated using the formula
@@
and the batch $\mathcal{B}_{k+1}$ is augmented with the above batch size. The constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ have to be chosen in practice. In implementation, we might also use the inner product test alone, since the (near) orthogonality of $\nabla F_{\mathcal{B}_{k}}(x_{k})$ and $\nabla F(x_{k})$ has not been observed in practice as mentioned in Bollapragada et al. (2018a), and the computation of the quantities in (10) would induce extra computational overhead for large-scale training. Then the batch size is updated using the formula
$b_{k+1}=\left\lceil\frac{\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)}{\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4}}\right\rceil.$
In practical implementation, similar to the norm test, we also rely on the heuristic approach which assumes the following three approximate equalities:
$\mathrm{Var}_{i\in\mathcal{B}_{k+1}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k+1}}(x_{k})\rangle)\approx\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle),$
$\mathrm{Var}_{i\in\mathcal{B}_{k+1}}\left(\nabla f(x_{k};\xi_{i})-\frac{\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k+1}}(x_{k})\rangle}{\|\nabla F_{\mathcal{B}_{k+1}}(x_{k})\|^{2}}\right)\approx\mathrm{Var}_{i\in\mathcal{B}_{k}}\left(\nabla f(x_{k};\xi_{i})-\frac{\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle}{\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}}\right),$
and $\|\nabla F_{\mathcal{B}_{k+1}}(x_{k})\|\approx\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|$, where $\mathcal{B}_{k}$ is the original batch and $\mathcal{B}_{k+1}$ is the new larger batch. These heuristics, despite their possible unreliability in theory, have worked well in experiments as mentioned in Bollapragada et al. (2018a).
@@ 4.3 Adaptive Sampling Methods for Adaptive Gradient Methods
with its approximate form being
$\frac{1}{b_{k}(b_{k}-1)}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\leqslant\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}.$
However, the existence of inner products in the inner product test and the orthogonality test precludes the derivation of their coordinate-specific versions.
@@ 5 Convergence Analysis
We provide two sets of convergence rate results for AdAdaGrad and AdAdaGrad-Norm. We first establish the convergence rate results of the norm test and the augmented inner product test for AdaGrad-Norm and AdaGrad for smooth nonconvex objectives. Then, we provide a convergence rate analysis of AdAdaGrad-Norm for non-uniformly smooth objectives. Proofs of this section can be found in Appendix B.
5.1 Assumptions
We might invoke the following assumptions in our analysis.
@@ 5.1 Assumptions
The function $F\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is continuously differentiable, bounded below by $F^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}F(x)=F(x^{\star})\in\mathbb{R}$ with $x^{\star}=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}F(x)$, and $L$-Lipschitz smooth, i.e., its gradient $\nabla F$ is $L$-Lipschitz continuous with constant $L\in\mathbb{R}_{+}$: for any $x,y\in\mathbb{R}^{d}$, $\|\nabla F(x)-\nabla F(y)\|\leqslant L\left\lVert x-y\right\rVert$.
The uniform smoothness condition often proves to be excessively strict for practical applications (compare this with a uniform upper bound on the spectrum of the Hessian of $F$ when $F$ is twice differentiable, denoted $\nabla^{2}F\preccurlyeq L\mathrm{Id}$). For example, Zhang et al. (2020a, b) have shown through numerical evidence that transformer architectures in language models exhibit loss landscapes that either do not meet the Lipschitz smoothness assumption or have significantly large Lipschitz constants $L$. To address such issues, Zhang et al. (2020b) propose the following more relaxed generalized smoothness condition:
@@
The function $F$ is $(L_{0},L_{1})$-smooth, i.e., for any $x,y\in\mathbb{R}^{d}$,
$\left\lVert\nabla F(x)-\nabla F(y)\right\rVert\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x)\right\rVert)\left\lVert x-y\right\rVert.$
@@
Assumption 3 (Expected strong growth).
For any batch $\mathcal{B}\subset\llbracket n\rrbracket$, the expected strong growth (E-SG) condition is given by
$(\forall x\in\mathbb{R}^{d})\quad\mathbb{E}[\|\nabla F_{\mathcal{B}}(x)\|^{2}]\leqslant\tau\|\nabla F(x)\|^{2},$
for some constant $\tau\in\mathbb{R}_{++}$. This condition can be relaxed in a sense that it only holds at each iteration, i.e.,
@@
for some constant $\tau\in\mathbb{R}_{++}$ and $(x_{k})_{k\in\mathbb{N}^{*}}$ are the iterates generated by some stochastic gradient methods.
@@
For every iteration $k\in\mathbb{N}^{*}$, if the conditions of the exact variance norm test hold with constant $\eta\in\left]0,1\right[$ and the conditions of the exact variance augmented inner product test hold with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ respectively, then the E-SG condition (12) holds with constant $\tau\in\mathbb{R}_{++}$, where $\tau=1+\eta^{2}$ for the exact variance norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the exact variance inner product test respectively.
A more precise statement and its proof are given in Appendix B. Although the expected strong growth condition holds in over-parameterized models (Vaswani et al., 2019) and is widely adopted in deep learning, the constant $\tau$ is usually problem-specific and unknown. In contrast, the two adaptive sampling methods impose a much weaker assumption; they require that this condition holds only for the iterates $(x_{k})_{k\in\mathbb{N}^{*}}$. In a practical implementation, the chosen constant $\tau$ is pre-specified and fixed, which is also likely to be smaller than the value for the global expected strong growth condition. If the (approximate) condition does not hold, the batch size is increased so that it holds for the next iterate (see Section 4).
@@ 5.2 Convergence Results
We first establish high probability convergence rates of the adaptive batch size schemes for AdaGrad-Norm and AdaGrad, substantially extending existing convergence rate results (in expectation) for SGD (e.g., De et al., 2016, 2017; Bollapragada et al., 2018a).
AdaGrad-Norm with the norm test or the augmented inner product test with any constant initial step size $\alpha\in\mathbb{R}_{++}$ enjoys the following high probability convergence bound for nonconvex functions.
@@
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\mathbb{N}^{*}}$ be the AdaGrad-Norm iterates (2) with any step size $\alpha\in\mathbb{R}_{++}$, where the batch sizes $(b_{k})_{k\in\mathbb{N}^{*}}$ are chosen such that either the (exact variance) norm test (6) with constant $\eta\in\left[0,1\right[$ or the (exact variance) augmented inner product test (7) and (9) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. Then, with probability at least $1-\delta$, we have
The constants $c_{1}$, $c_{2}$ and $c_{3}$ are defined through $c_{1}\coloneqq\frac{2}{\alpha(1-\rho^{-1})}(F(x_{1})-F^{\star}+\tau\alpha\left\lVert\nabla F(x_{0})\right\rVert^{2}/(2\sqrt{v_{0}})+\tau L^{2}\alpha^{3}(1+\rho\tau)/\sqrt{v_{0}}-L\alpha^{2}\log(v_{0})/2)$, $c_{2}\coloneqq 2L\alpha/(1-\rho^{-1})$, and $c_{3}\coloneqq 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)$, where $\rho\in\left]1,+\infty\right[$ is a constant, with $\tau=1+\eta^{2}$ for the norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the augmented inner product test.
@@
For every iteration $k\in\mathbb{N}^{*}$, if the condition of the coordinate-wise exact variance norm test (11) holds with constant $\eta\in\left]0,1\right[$, then the coordinate-wise E-SG condition is satisfied at each iteration $k\in\mathbb{N}^{*}$, i.e., for all $(k,j)\in\mathbb{N}^{*}\times\llbracket d\rrbracket$, we have
$\mathbb{E}_{k}\left[\left(\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\right]\leqslant\tau(\partial_{j}F(x_{k}))^{2},$ (13)
where $\tau=1+\eta^{2}$.
@@
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\mathbb{N}^{*}}$ be the AdaGrad iterates (1) with any step size $\alpha\in\mathbb{R}_{++}$, where the batch sizes $(b_{k})_{k\in\llbracket K\rrbracket}$ are chosen so that the coordinate-wise (exact variance) norm test (11) is satisfied at each iteration. Then, with probability at least $1-\delta$, we have
$\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(\frac{1}{K\delta^{2}}\right).$
The above theorem establishes a sublinear convergence rate (with high probability) for AdAdaGrad for nonconvex functions, while such a rate (in expectation) has been established for SGD in Bollapragada et al. (2018a, Theorem 3.4).
@@
Suppose that Assumption 2 holds. Let $(x_{k})_{k\in\mathbb{N}^{*}}$ be the AdaGrad-Norm iterates (2). Also suppose that the batch sizes $(b_{k})_{k\in\mathbb{N}^{*}}$ are chosen so that either the (exact variance) norm test (6) with constant $\eta\in\left[0,1\right[$ or the (exact variance) augmented inner product test (7) and (9) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. If the AdaGrad-Norm step size $\alpha\leqslant\min\{\omega/(2\rho_{1}\tau),\sqrt{\omega/(2\rho_{1}\tau)}\}/L_{1}$, where $(\rho_{1},\rho_{2},\omega)\in\mathbb{R}_{++}^{3}$ are constants such that $11/\rho_{1}+\rho_{1}/\rho_{2}+2\omega<$ 1 and $\tau\geqslant\min\{\omega/(2\min\{\rho_{1},\rho_{2}\}),1\}$, with $\tau=1+\eta^{2}$ for the norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the augmented inner product test, then with probability at least $1-\delta$, we have
$\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(\frac{1}{K\delta^{2}}\right).$

### v2

@@ 4 Adaptive Sampling Methods
The adaptive batch schemes proposed in this paper are based on a family of adaptive sampling methods for stochastic optimization problems. Although such methods are extended to problems with different structures (see Section 2 for details), here we focus on the simplest unconstrained setting.
@@ 4.1 Norm Test
Byrd et al. (2012) proposed an adaptive sampling method, called the norm test or the norm condition, that was derived according to the following fundamental fact in stochastic optimization. If $f(\cdot;\xi)$ is a convex and continuously differentiable function for any $\xi\in\mathcal{Z}$, then $-\nabla F_{\mathcal{B}}(x)$ is a descent direction for $F$ at $x\in\mathbb{R}^{d}$ if there exists $\eta\in\left[0,1\right)$ such that
@@
See Appendix C for its proof. When the sample size $n$ is large, $\nabla F$ is not available, but $\delta_{\mathcal{B}}(x)$ can be approximated by
@@
Therefore, as the sample size $n\to\infty$, condition (1) can be approximated as $\frac{1}{b}\mathrm{Var}_{i\in\mathcal{B}}(\nabla f(x;\xi_{i}))\leqslant\eta^{2}\|\nabla F_{\mathcal{B}}(x)\|^{2}$, which also corresponds to the case of sampling with replacement. In a practical implementation, the dynamic sample gradient algorithm with the norm test (Byrd et al., 2012) performs the following SGD update: $x_{k+1}=x_{k}-\alpha_{k}\nabla F_{\mathcal{B}_{k}}(x_{k})$ with step size $\alpha_{k}>0$, and then checks the following (approximate) norm test condition:
@@
meaning that the condition (1) holds in expectation. Here we abbreviate the conditional expectation on $\mathcal{F}_{k}$ (that is, the sigma algebra $\mathcal{F}_{k}\coloneqq\sigma(\{x_{1},\mathcal{B}_{1},\ldots,\mathcal{B}_{k-1}\})$) by $\mathbb{E}_{k}[\cdot]\coloneqq\mathbb{E}[\cdot\,|\,\mathcal{F}_{k}]$. Should the condition (3) not be met, the subsequent batch size is determined by the formula: $b_{k+1}=\left\lceil\mathrm{Var}_{i\in\mathcal{B}_{k}}(\nabla f(x_{k};\xi_{i}))/\left(\eta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{2}\right)\right\rceil$, and the batch $\mathcal{B}_{k+1}$ expands accordingly. The selection of the constant $\eta$ influences the probability of obtaining a descent direction, affecting the probability of increasing the batch size.
@@ 4.2 Inner Product Test
Bollapragada et al. (2018a) noted that the norm test often leads to a rapid increase in batch size, which could negate the benefits of adaptive sampling, and proposed the inner product test as an alternative. This test further controls the variance of the term $\langle\nabla F_{\mathcal{B}_{k}}(x_{k}),\nabla F(x_{k})\rangle$ to ensure that $-\nabla F_{\mathcal{B}_{k}}(x_{k})$ is a descent direction with high probability. The test selects a large enough batch size $b_{k}$ to meet the (exact variance) inner product test condition, which is satisfied when there is a $\vartheta>0$ ensuring that:
@@
where the variance on the left hand side can be evaluated by (2). To establish a convergence rate, it is crucial to eliminate the possibility that $\nabla F_{\mathcal{B}_{k}}(x_{k})$ and $\nabla F(x_{k})$ are (near) orthogonal. Motivated by this need for theoretical guarantees, Bollapragada et al. (2018a) also introduce the following condition of the (exact variance) orthogonality test: there exists $\nu>0$ such that
@@
The above exact test also takes an approximate form for implementation purposes, given by
@@
Note that both (6) and (8) have to be satisfied, which collectively leads to the (approximate) augmented inner product test. If any of the conditions are violated, the batch size will be updated using the formula
@@
and the batch $\mathcal{B}_{k+1}$ is augmented with the above batch size. The constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ have to be chosen in practice. In implementation, we might also use the inner product test alone, since the (near) orthogonality of $\nabla F_{\mathcal{B}_{k}}(x_{k})$ and $\nabla F(x_{k})$ has not been observed in practice as mentioned in Bollapragada et al. (2018a), and the computation of the quantities in (8) would induce extra computational overhead for large-scale training. Then the batch size is updated using the formula $b_{k+1}=\left\lceil\mathrm{Var}_{i\in\mathcal{B}_{k}}(\langle\nabla f(x_{k};\xi_{i}),\nabla F_{\mathcal{B}_{k}}(x_{k})\rangle)/\left(\vartheta^{2}\|\nabla F_{\mathcal{B}_{k}}(x_{k})\|^{4}\right)\right\rceil$.
@@ 4.3 Adaptive Sampling Methods for Adaptive Gradient Methods
We focus on two simple adaptive gradient methods, AdaGrad (Duchi et al., 2011; McMahan and Streeter, 2010) and AdaGrad-Norm (Streeter and McMahan, 2010), whose step sizes are computed adaptively based on the magnitudes of previous (batch) stochastic gradients. AdaGrad was proposed for online convex optimization, which takes the following update:
$(\forall k\in\mathbb{N}^{*})\quad v_{k}=v_{k-1}+g_{k}^{2},\quad x_{k+1}=x_{k}-\alpha g_{k}\odot v_{k}^{-\nicefrac{{1}}{{2}}},$ (9)
where $g_{k}\coloneqq\nabla F_{\mathcal{B}_{k}}(x_{k})$, $\alpha>0$ is a constant step size (or learning rate), $\odot$ denotes the Hadamard product, and the power operations are performed coordinate-wise. Since $(v_{k})_{k\in\mathbb{N}}\subset\mathbb{R}_{++}^{d}$ is a sequence of positive vectors, AdaGrad has adaptive coordinate-wise step sizes.
$(\forall k\in\mathbb{N}^{*})\quad v_{k}=v_{k-1}+\left\lVert g_{k}\right\rVert^{2},\quad x_{k+1}=x_{k}-\alpha g_{k}/\sqrt{v_{k}},$ (10)
@@
with its approximate form being $\frac{1}{b_{k}(b_{k}-1)}\sum_{i\in\mathcal{B}_{k}}\left(\partial_{j}f(x_{k};\xi_{i})-\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\leqslant\eta^{2}(\partial_{j}F_{\mathcal{B}_{k}}(x_{k}))^{2}$. However, the existence of inner products in the inner product test and the orthogonality test precludes the derivation of their coordinate-specific versions.
@@ 5 Convergence Analysis
We provide two sets of convergence rate results for AdAdaGrad and AdAdaGrad-Norm. We first establish the convergence rate results of the norm test and the augmented inner product test for AdaGrad-Norm and AdaGrad for smooth nonconvex objectives. Then, we provide a convergence rate analysis of AdAdaGrad-Norm for non-uniformly smooth objectives. Proofs of this section can be found in Appendix C. We might invoke the following assumptions in our analysis.
@@
The function $F\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is continuously differentiable, bounded below by $F^{\star}\coloneqq\inf_{x\in\mathbb{R}^{d}}F(x)=F(x^{\star})\in\mathbb{R}$ with $x^{\star}=\operatorname*{argmin}_{x\in\mathbb{R}^{d}}F(x)$, and $L$-Lipschitz smooth, i.e., its gradient $\nabla F$ is $L$-Lipschitz continuous with constant $L\geqslant 0$: for any $x,y\in\mathbb{R}^{d}$, $\|\nabla F(x)-\nabla F(y)\|\leqslant L\left\lVert x-y\right\rVert$.
[…]
@@
The function $F$ is $(L_{0},L_{1})$-smooth with constants $(L_{0},L_{1})\in\mathbb{R}_{+}^{2}$, i.e., for any $x,y\in\mathbb{R}^{d}$, $\left\lVert\nabla F(x)-\nabla F(y)\right\rVert\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x)\right\rVert)\left\lVert x-y\right\rVert$.
@@
Definition 1 (Expected strong growth).
For any batch $\mathcal{B}\subset\llbracket n\rrbracket$, the expected strong growth (E-SG) condition is given by $(\forall x\in\mathbb{R}^{d})\;\mathbb{E}[\|\nabla F_{\mathcal{B}}(x)\|^{2}]\leqslant\tau\|\nabla F(x)\|^{2}$, for some constant $\tau>0$. This condition can be relaxed in a sense that it only holds at each iteration, i.e.,
@@
for some constant $\tau>0$ and $(x_{k})_{k\in\mathbb{N}^{*}}$ are the iterates generated by some stochastic gradient methods.
@@
For every iteration $k\in\mathbb{N}^{*}$, if the conditions of the exact variance norm test hold with constant $\eta\in\left(0,1\right)$ and the conditions of the exact variance augmented inner product test hold with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ respectively, then the E-SG condition (12) holds with constant $\tau>0$, where $\tau=1+\eta^{2}$ for the exact variance norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the exact variance inner product test respectively.
A more precise statement and its proof are given in Appendix C. Although the expected strong growth condition holds in over-parameterized models (Vaswani et al., 2019) and is widely adopted in deep learning, the constant $\tau$ is usually problem-specific and unknown. In contrast, the two adaptive sampling methods impose a much weaker assumption; they require that this condition holds only for the iterates $(x_{k})_{k\in\mathbb{N}^{*}}$. In a practical implementation, the chosen constant $\tau$ is pre-specified and fixed, which is also likely to be smaller than the value for the global expected strong growth condition. If the (approximate) condition does not hold, the batch size is increased so that it holds for the next iterate (see Section 4).
@@ 5.1 Convergence Results
We first establish high probability convergence rates of the adaptive batch size schemes for AdaGrad-Norm and AdaGrad, substantially extending existing convergence rate results (in expectation) for SGD; see e.g., De et al. (2016, 2017); Bollapragada et al. (2018a). AdaGrad-Norm with the norm test or the augmented inner product test with any constant initial step size $\alpha>0$ enjoys the following high probability convergence bound for nonconvex functions.
@@
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\llbracket K\rrbracket}$ be the AdaGrad-Norm iterates (10) with any step size $\alpha>0$, where the batch sizes $(b_{k})_{k\in\llbracket K\rrbracket}$ are chosen such that either the (exact variance) norm test (4) with constant $\eta\in(0,1)$ or the (exact variance) augmented inner product test (5) and (7) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. Then, with probability at least $1-\delta$, we have $\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(1/(K\delta^{2})\right)$.
@@
For every iteration $k\in\mathbb{N}^{*}$, if the condition of the coordinate-wise exact variance norm test (11) holds with constant $\eta\in(0,1)$, then the coordinate-wise E-SG condition is satisfied at each iteration $k\in\mathbb{N}^{*}$, i.e., for all $(k,j)\in\mathbb{N}^{*}\times\llbracket d\rrbracket$, we have
$\mathbb{E}_{k}\left[\left(\partial_{j}F_{\mathcal{B}_{k}}(x_{k})\right)^{2}\right]\leqslant(1+\eta^{2})(\partial_{j}F(x_{k}))^{2}.$ (13)
@@
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\llbracket K\rrbracket}$ be the AdaGrad iterates (9) with any step size $\alpha>0$, where the batch sizes $(b_{k})_{k\in\llbracket K\rrbracket}$ are chosen so that the coordinate-wise (exact variance) norm test (11) is satisfied at each iteration. Then, with probability at least $1-\delta$, we have $\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(1/(K\delta^{2})\right)$.
The above theorem establishes a sublinear convergence rate (with high probability) for AdAdaGrad for nonconvex functions, while such a rate (in expectation) has been established for SGD in Bollapragada et al. (2018a), Theorem 3.4.
@@
Suppose that Assumption 2 holds. Let $(x_{k})_{k\in\llbracket K\rrbracket}$ be the AdaGrad-Norm iterates (10). Also suppose that the batch sizes $(b_{k})_{k\in\llbracket K\rrbracket}$ are chosen so that either the (exact variance) norm test (4) with constant $\eta\in(0,1)$ or the (exact variance) augmented inner product test (5) and (7) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. If the AdaGrad-Norm step size $\alpha\leqslant\min\{\omega/(2\rho_{1}\tau),\sqrt{\omega/(2\rho_{1}\tau)}\}/L_{1}$, where $(\rho_{1},\rho_{2},\omega)\in\mathbb{R}_{++}^{3}$ are constants such that $11/\rho_{1}+\rho_{1}/\rho_{2}+2\omega<$ 1 and $\tau\geqslant\min\{\omega/(2\min\{\rho_{1},\rho_{2}\}),1\}$, with $\tau=1+\eta^{2}$ for the norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the augmented inner product test, then with probability at least $1-\delta$, we have $\min_{k\in\llbracket K\rrbracket}\left\lVert\nabla F(x_{k})\right\rVert^{2}=\mathscr{O}\left(1/(K\delta^{2})\right)$.